OpenStudy (anonymous):

fraction division 1/(2/3) and (2/3)/2

OpenStudy (anonymous):

Fraction division is the inverse of fraction multiplication. 1 / (2/3) = (1/1) * (2/3) = (2/3) (2/3) / 2 = (2/3) * (1 / 2) = (2/6) = (1/3)

OpenStudy (anonymous):

I thought (1/1)*(3/2)?? for the first one

OpenStudy (anonymous):

dont u swap the fraction

OpenStudy (anonymous):

mmm

OpenStudy (anonymous):

yes, you "swap" the fraction. \[\frac{a}{b} \div \frac{c}{d} = \frac{b}{a} * \frac{c}{d}\]

OpenStudy (anonymous):

so why the frist answer is 2/3? if its swamp then it should be 3/2

OpenStudy (anonymous):

?